# Dyons in higher-dimensional gauge theories

### v1

@@ 6 Summary and discussion
As a remarkable specific feature of this type of theories, especially the 5D GHU compactified on a circle, the Higgs potential is periodic in the Higgs field, reflecting the topological nature of the extra dimension. This property was argued to play central roles in deriving the multiple novel predictions made by the higher dimensional gauge theory, demonstrated in this paper.
@@
Next, the argument was generalized to the case of dyon. It turns out that one of the BPS conditions still can be written as an (anti-)self-dual-like condition, similar to that in the case of BPS monopole, and the mass of the BPS dyon, $M_{{\rm BPS}}$ is still proportional to the quantized Higgs VEV in the context of the higher dimensional gauge theory. $M_{{\rm BPS}}$, however, also depends on a parameter $\mu$, denoting the ratio of the electric and magnetic charges of the dyon, which is usually left as a free parameter.
In this paper, we proposed a novel mechanism to quantize the parameter $\mu$, again utilizing the characteristic features of the higher dimensional gauge theories. Namely, we show that in the 5D gauge theories, the CS term is induced at the quantum level, and replacing the extra space component of the introduced U(1) gauge field by its VEV, the $\theta$ term arises. Importantly , here again the VEV is quantized and through the Witten effect we have reached to a remarkable conclusion that the electric charge of the dyon, and therefore the parameter $\mu$ is discretized. This means that $M_{{\rm BPS}}$ is also discretized.

### v2

@@ 6 Summary and discussion
As a specific feature of this type of theories, especially the 5D GHU compactified on a circle, the Higgs potential is periodic in the Higgs field, reflecting the topological nature of the extra dimension. This property was argued to play central roles in deriving the multiple characteristic predictions made by the higher dimensional gauge theory, demonstrated in this paper.
@@
As was discussed in some details in section 2, from the viewpoint of the caloron [19, 20, 21], which may be understood to be a composition of a pair of constituent monopole and anti-monopole, our conclusion of the quantized monopole mass corresponds to the specific case, where only a single monopole exists that carries the whole topological charge.
Next, the argument was generalized to the case of dyon. It turns out that one of the BPS conditions still can be written as an (anti-)self-dual-like condition, similar to that in the case of BPS monopole, and the mass of the BPS dyon, $M_{{\rm BPS}}$ is still proportional to the quantized Higgs VEV in the higher dimensional gauge theory. $M_{{\rm BPS}}$, however, also depends on a parameter $\mu$, denoting the ratio of the electric and magnetic charges of the dyon, which is usually left as a free parameter.
In this paper, we proposed an interesting mechanism to quantize the parameter $\mu$, again utilizing the characteristic features of the higher dimensional gauge theories. Namely, we showed that in the 5D gauge theories, the CS term is induced at the quantum level, and replacing the extra space component of the introduced U(1) gauge field by its VEV, the $\theta$ term arises. Importantly, here again the VEV is quantized and through the Witten effect we have reached to a conclusion that the electric charge of the dyon, and therefore the parameter $\mu$ is discretized. This means that $M_{{\rm BPS}}$ is also discretized.